For the month of April.

Costs:

  Work in process, April 1

    Direct materials: 100% complete

$100,000

    Conversion costs: 20% complete

70,000

      Cost of work in process, April 1

$170,000

Costs incurred during production in April

    Direct materials

$ 800,000

    Conversion costs

365,000

      Costs incurred in April

$1,165,000


Units transferred out totaled 17,000. Ending work in process was 1,000 units that are 100% complete as to materials and 40% complete as to conversion costs.

Using the results of the equivalents units found in FIB 1, compute unit cost per

(a) material

(b) conversion cost
Answer #1

Pgo Answer @ compute unit cost per material: Equivalent unit de material = 17000+ (1000x1004.) = 17000+ 1000 = 18,000 units I